OpenStudy (anonymous):

a triangle has side lenths of 9cm, 40cm, 41cm, show mathematicly that it is a right angled triangle

OpenStudy (anonymous):

do u know pythagores theorem ??????????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just apply that

OpenStudy (anonymous):

but i dont know how to explain

OpenStudy (unklerhaukus):

the hypotenuse it the longest side of the triangle

OpenStudy (anonymous):

prove lhs = rhs H^2=P^2+B^2

OpenStudy (anonymous):

where H is hypotenuse P prpendicular B base

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

what is H here?

OpenStudy (anonymous):

41 cm

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then H^2=?

OpenStudy (anonymous):

whats the triangle thing

OpenStudy (anonymous):

its representation of square

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

since there is no small subscript

OpenStudy (anonymous):

@UnkleRhaukus how to show squre of a no?

OpenStudy (unklerhaukus):

``` \[H^2=P^2+B^2\] ``` \[H^2=P^2+B^2\]

OpenStudy (unklerhaukus):

\[41^2=41\times41\\=(40+1)\times(40+1)\\=40(40+1)+1(40+1)\\=\]

OpenStudy (anonymous):

ohk u r genius :)

OpenStudy (anonymous):

a^2 + b^2 = c^2 (c is the hypotenuse) 9^2 + 40^2 = 41^2 81 + 1600 = 1681 1681 = 1681 (correct)

OpenStudy (unklerhaukus):

40(40+1)+1(40+1) =1600+40+40+1 =1681